# XOXO: Stealthy Cross-Origin Context Poisoning Attacks against AI Coding Assistants

### v1

@@ 5 Greedy Cayley Graph Search
The goal of XOXO attack is to modify the input code through semantics-preserving adversarial transformations that deceive the LLM, without changing the code’s underlying logic. Simple transformations include renaming variables, switching between equivalent control structures (e.g., a for loop to a while loop), or reordering independent statements. Although these transformations preserve semantics, they still can change model output and confidence, as also shown by prior works [25, 26]. These simple, atomic transformations can be combined to form composite transformations that also preserve semantics, creating a vast space of potential transformations. The attack must, therefore, efficiently explore this space to identify transformations that induce incorrect model outputs.
We observe that applying multiple confidence-reducing transformations often has a monotonic effect i.e. as more of these transformations are composed, the model’s confidence decreases consistently. We test this hypothesis in Appendix C. Our search algorithm, GCGS, leverages this monotonicity to efficiently find a transformation composition that lowers model confidence below a threshold, resulting in incorrect output. We observe that the transformation space forms a tree structure, where each node (a composite transformation) is generated from a sequence of atomic transformations, and all nodes at a given level in the tree contain the same number of transformations. GCGS leverages this tree structure by representing it as a Cayley Graph [27], with the attack strategy executing a greedy walk through the graph to exploit the monotonic decrease in model confidence along the path.
@@ 5.1 Preliminaries
We consider a generating set $G$ of atomic transformations that generates the entire group of complex transformations. Each transformation $g_{i}\in G$ maps a code snippet $\mathcal{C}$ to $\mathcal{C}^{\prime}$ through atomic changes, such as replacing every occurrence of an identifier foo with bar, while preserving code semantics. For each transformation $g_{i}$, there exists an inverse transformation $g_{i}^{-1}\in G^{-1}$ that reverses its effect (e.g., replacing bar back to foo), such that their composition yields an identity transformation $g_{\textstyle i}\cdot g_{i}^{-1}=e$.
Since there are no restrictions on composing transformations in $G$, this set forms a free group $F(G)$ generated by $G$, where each element represents a sequence of transformations from $G\cup G^{-1}$. To visualize and systematically explore potential transformation sequences, we can represent this group structure using a Cayley Graph. While Cayley Graphs can be used to represent any group, in the case of a free group, this graph becomes an infinite tree $\mathcal{T}$ as shown in Figure 3. In $\mathcal{T}$, each vertex represents an element of $F(G)$ (a composite transformation), and each edge represents the application of a single transformation $g\in(G\cup G^{-1})\setminus{e}$. Unlike other tree structures, Cayley graphs naturally handle cases where different transformation sequences when composed produce identical code snippets.
@@ 5.2 Problem Formulation
Now consider a code model $\mathcal{M}:\ \mathcal{C}\rightarrow\mathcal{Y}$, mapping code snippets to an output space $\mathcal{Y}$ (e.g., class labels for classification tasks or token sequences for generation tasks). For many downstream tasks, even with black-box access to $\mathcal{M}$, we can approximately measure the model’s confidence in its predictions. Let $\alpha:\mathcal{C}\rightarrow[0,1]$ be a confidence scoring function. For classification tasks, $\alpha(c)$ can be derived from the probability distribution over classes [28, 29]. For generation-based tasks with current LLMs, we can compute $\alpha(c)$ using perplexity or the prediction stability when sampling with moderate temperature. This provides us with a continuous measure of the model’s certainty in its predictions, where lower values of $\alpha(g_{i}(c))$ indicate that applying transformation $g_{i}$ makes the model less confident about its output.
The model’s predictions are impacted by the program semantics-preserving transformations, indicating that the model does not completely capture the semantic meaning and instead partially relies on spurious patterns in its decision-making process. This reliance on spurious patterns suggests that accumulating multiple semantic-preserving changes will progressively degrade the model’s understanding of the program, manifesting as decreased confidence in its predictions. For transformations $g_{i},g_{j}\in G$, we expect their composition to result in lower confidence: $\alpha(g_{i}\cdot g_{j}(c))<\min(\alpha(g_{i}(c)),\alpha(g_{j}(c)))$. We test this theoretical intuition in Appendix C, where we find that composing transformations that individually decrease model confidence indeed yields compounded decreases in confidence, exhibiting monotonicity with respect to path length in $\mathcal{T}$. This monotonic relationship between path length and confidence scores suggests a greedy strategy for finding adversarial transformations. By following paths of decreasing confidence in $\mathcal{T}$, we can efficiently reach composite transformations which lead the model to mispredict.
@@ 5.3 GCGS Algorithm
Leveraging the monotonicity property, GCGS efficiently finds a path to a transformation $\tilde{g}$ such that $\mathcal{M}(\tilde{g}(c))\neq\mathcal{M}(c)$. As illustrated in Figure 3, the algorithm explores the Cayley Graph $\mathcal{T}$ in two phases (Algorithm 1). The first phase is Shallow Exploration, where GCGS begins by sampling a set $G^{R}\subset(G\cup G^{-1})\setminus{e}$ of $R$ generators. For each $g\in G^{R}$, it computes and stores the model confidence $\alpha(g(c))$ in a $g$-$\alpha$ map $A$. If any of these atomic transformations directly causes a model failure, the transformed code snippet is returned as the solution. In the next phase, Deep Greedy Composition, which is when no atomic transformation succeeds, GCGS leverages the stored confidence values to greedily compose transformations. Starting with the identity transformation $\tilde{g}=e$, it iteratively composes $\tilde{g}$ with generators from $G^{R}$, considering them in order of increasing confidence values in $A$. This effectively implements a guided descent through $\mathcal{T}$ towards likely failure points. Moreover, the inverse transformations in the generating set ($G^{-1}$) give GCGS, an additional capability to revert any applied transformation along the greedy walk. GCGS repeats these two phases, maintaining the confidence map $A$ across iterations until it either finds an adversarial example or reaches the maximum query limit.
5.4 GCGS with Warm-up
In the shallow exploration phase of the GCGS, the random sampling from $(G\cup G^{-1})\setminus{e}$ in each iteration to form $G^{R}$ can be query-inefficient as the sample may contain fewer confidence-reducing transformations. In practice, certain transformations might consistently be more effective at reducing model confidence across similar code snippets. We can exploit this pattern to make GCGS more efficient.
Consider an attacker with access to code snippets $C^{W}\sim C$ drawn from the same distribution as the target snippets. We use $C^{W}$ in an offline stage to learn which transformations tend to be most effective, warming up our attack to sample $G^{R}$ more intelligently during the shallow exploration. We split $C^{W}$ into training set $C^{T}$ and validation set $C^{V}$. Over multiple rounds, we randomly sample $G^{R}$ from $(G\cup G^{-1})\setminus{e}$ and record $\alpha(g(c))$ for each $g\in G^{R}$ and $c\in C^{T}$. Using the average confidence drop of each transformation in $G^{R}$ on $C^{T}$, we run GCGS on $C^{V}$ to validate if the current sample of $G^{R}$ is better than the previous round. The warm-up procedure keeps refining the set $G^{R}$ until either we saturate, with GCGS’s performance on $C^{V}$ starting to deteriorate or the maximum number of rounds is reached.
We note that the warm-up phase is an offline process conducted without the test code snippets. As we show in § 7.3, this phase is optional but boosts GCGS’s efficiency during the attack stage. Furthermore, we reveal in § 7.4 that the warm-up phase transfers across different model architectures, suggesting that certain code transformations are consistently more likely to expose model vulnerabilities. Based on these findings, we believe there is a potential for a universal attack [30], an exciting direction for future work.
Algorithm 1 Greedy Cayley Graph Search
@@ 5.4 GCGS with Warm-up
6 Implementation
Transformations. Although the Cayley Graph structure accommodates any semantics-preserving transformations (including non-commutative ones), for attack implementation we focus on identifier replacements, specifically function, parameter, variable, and class-member names. This is because identifier replacements offers a larger search space compared to other transformations like control flow modifications, while enabling precise atomic control over the magnitude of code changes.
We leverage tree-sitter [31] to parse code snippets and extract identifier positions. To maintain natural and realistic transformations, we employ the different identifier sourcing strategies for each task. For defect and clone detection tasks, we seed identifiers from their respective training sets to avoid out-of-distribution effects in fine-tuned models. For smaller Python datasets (HumanEval+ and MBPP+), we extract identifiers from CodeSearchNet/Python [32] to ensure sufficient variety. HumanEval+ and MBPP+ tasks additionally incorporate Python input-output assertions in docstrings (e.g., $>$ string_xor(’010’, ’110’) ’100’ or assert is_not_prime(2) == False), we maintain consistency by replacing function names in both the code and assertions as done by previous implementations  [25, 26]. This consistency is crucial as the assertions are part of the model’s input, and any naming discrepancies would test the model’s ability to handle inconsistent references rather than its semantic understanding.
When constructing composite transformations, we iterate through identifier-replacement pairs ordered by increasing model confidence (based on the stored $g$-$\alpha$ map). For classification tasks, we measure the model’s confidence as the probability of predicting the correct class. For generation tasks, we measure the model’s confidence as the sum of the generated tokens’ log probabilities. At each iteration we select the highest-confidence pair where neither the identifier nor its replacement appears in previous steps. This process continues until we either discover a breaking transformation or exhaust maximum number of queries to the model.
Warm Up. To highlight the practicality of attack warm-up, we use a small 5% sample of the model’s training and validation datasets for reasoning tasks, illustrating that an attacker requires minimal access to in-distribution examples for effective results. This set ($C^{W}$) is kept disjoint from the model’s fine-tuning set to ensure fair evaluation. For code completion tasks, where dataset sizes are smaller, we implement cross-dataset warm-up (e.g., using MBPP+ to initialize attacks on HumanEval+ and vice versa). The warm-up process begins with randomly sampling replacements for each identifier in the training set code snippets ($C^{T}$) and tracking the average drop in model confidence for each replacement across the complete $C^{T}$. Based on the top performing replacements from $C^{T}$, an attack is executed on $C^{V}$ for getting each replacement’s validation performance score. Using this score, we select top-$k$ highest-scoring transformations as warm-up set for the actual attack. We also experimented with alternative sampling methods, including distribution biasing and softmax-based sampling, but found that the straightforward top-$k$ selection strategy provided the best results.
@@ 7 Evaluation
The evaluation addresses the following research questions:
  • RQ1: Is GCGS effective at attacking SoTA LLMs on in-context code generation?
  • RQ2: Can GCGS make SoTA LLMs inject vulnerabilities while preserving functional correctness?
  • RQ3: Is GCGS effective at attacking SoTA LLMs on code reasoning?
  • RQ4: How effectively can attack knowledge be transferred between models through cross-model hot start?
  • RQ5: Can GCGS be used to inject subtle bugs?
  • RQ6: Can adversarial fine-tuning mitigate GCGS?
@@
To answer RQ1, we devise a new in-context code generation task based on industry-standard benchmarks that simulate how a typical AI assistant would formulate code generation prompts with supplementary code context. We then evaluate how vulnerable the underlying LLMs are to context poisoning through semantics-preserving code transformations. For RQ2, we evaluate the perplexity-guided GCGS (GCGS+P) on the CWEval dataset, a benchmark designed to assess the security and functional correctness of LLM-generated code. For RQ3, we compare against SoTA adversarial attacks on two security-critical code reasoning tasks: defect detection and clone detection. To answer RQ4, we compare GCGS against its variant warmed up on a surrogate model (GCGS+W) as described in § 5.4. For RQ5, we investigate the foundational ability of GCGS to inject specific bugs by using its capability to cause failures in a small set of test cases for a code generation task as a proxy signal. Finally, for RQ6, we conduct adversarial fine-tuning on target models using examples generated by GCGS and subsequently evaluate whether GCGS can discover new adversarial examples on these fine-tuned models.
@@
Model Confidence (Generation vs Reasoning) In code reasoning tasks involving classification, adversarial attacks can use the model’s output probabilities as a direct optimization signal. These approaches leverage the probability distribution over classes as a fitness function, guiding input transformations to reduce the model’s confidence in the correct class. Our GCGS is generic and can employ this guidance. For code models where this guidance exists, we compare GCGS and GCGS with warm-up (GCGS+W) against SoTA black-box adversarial input generation methods (ALERT [28], MHM [33], RNNS [29], and WIR-Random [34]), which use model confidence to guide their exploration.
However, existing black-box approaches fundamentally fail when attacking models on code generation. While token-level probabilities or aggregate metrics like perplexity are accessible, they do not directly reflect the model’s confidence or output quality, making them unreliable signals for guiding attacks. Furthermore, SoTA adversarial methods designed for code reasoning tasks become computationally infeasible for generative tasks, as each evaluation requires generating a full sequence rather than a single classification decision. Therefore, for generative tasks, we cannot compare GCGS to the same methods we compared against in code reasoning.
In the generation setting, we evaluate GCGS both with perplexity feedback from generated inputs (GCGS+P) and without it when token probabilities are unavailable (GCGS). We assess all open-source models with and without perplexity feedback. For Claude 3.5 Sonnet, we evaluate GCGS without perplexity due to their API restrictions on accessing token-level log probabilities. In the case of GPT-4o, we solely evaluate GCGS+P with perplexity due to budget constraints and fewer queries made by GCGS when using perplexity as feedback. Due to cost considerations with closed-source models and limited compute resources for open-source models, we exclude GCGS with warm-up (GCGS+W) in the generation setting.
@@
Evaluation Metrics. We evaluate our attack using two categories of metrics: those measuring performance and those measuring the quality of generated adversarial examples. The performance of our attack is measured using two metrics: (i) Attack Success Rate (ASR) is the percentage of cases where an attack transforms correct model outputs into incorrect ones. This applies to classification (model changes from correct to incorrect class) and code completion (generated code changes from passing to failing test cases), and (ii) Number of Queries (# Queries) refers to the mean queries to the model per attack, indicating the attack’s efficiency under real-world constraints like rate limits and costs.
We evaluate the quality and naturalness of adversarial examples using three metrics widely adopted in prior work [28, 29, 35]: (i) CodeBLEU [36] measures code similarity by combining BLEU score with syntax tree and data flow matching, ranging from 0 (completely distinct) to 100 (identical). Higher scores indicate adversarial code that better preserves the original code’s structure and functionality, (ii), and (iii): Identifier and Position Metrics (# Identifiers, # Positions) count the number of replaced identifiers and their occurrences in the code. For instance, changing one variable used multiple times affects several positions. Lower numbers indicate more natural modifications that are harder to detect through static analysis or code review.
Machine Details. We conducted model fine-tuning and generative tasks using consumer hardware: a 20-core processor with 64GB RAM and dual NVIDIA RTX 3090 GPUs, running Ubuntu 22.04 and CUDA 12.1. For comparative evaluations against SoTA attacks, model transferability and adversarial fine-tuning experiments, we utilized GCP g2-standard-96 instances equipped with 96 cores, 384GB RAM, and eight NVIDIA L4 GPUs (one GPU per attack) on Debian 11 with CUDA 12.1. To serve LLMs, we use either transformers 4.42.4 [37] or vllm 0.6.3.post1 [38]. We access GPT-4o through OpenAI API and Claude 3.5 Sonnet through GCP Vertex AI API.
7.1 RQ1: In-Context Code Generation
@@
Task Description. We formulate a new task to simulate a code generation task with benign, AI Coding Assistant-supplied context. For code generation, we use the industry-standard HumanEval+ and MBPP+ datasets from EvalPlus [39, 40, 41]. HumanEval+ (164 problems) and MBPP+ (378 problems) feature more comprehensive test cases than HumanEval [40] and MBPP [41] to more accurately evaluate code generation performance. Both datasets contain Python functions with natural language descriptions and input-output examples, with performance measured using the pass@1 metric, meaning percentage of correct solutions a model generates on its first attempt. We selected these datasets because they are frequently used in model development and LLMs achieve high-performance on them, making them a conservative choice for demonstrating adversarial vulnerability – success on these familiar, well-mastered tasks would indicate fundamental weaknesses even in optimal conditions.
@@
We augment each target problem with three randomly sampled, solved examples from the same dataset to simulate a realistic code completion scenario for a developer using an AI assistant. The prompt instructs models to complete the target example while maintaining consistency with the coding style and naming conventions of the context examples. The complete prompt template is provided in Appendix A.
To evaluate our attack’s effectiveness, we target both closed-source (GPT-4o [42], Claude 3.5 Sonnet v2 (2024/10/22) [43]) and open-source models (Llama 3 8B Instruct [44], CodeQwen 1.5 7B Chat [45], DeepSeek Coder 6.7B Instruct [46], and Artigenz Coder DS 6.7B [47]). Several of these models, notably GPT-4o and Claude 3.5 Sonnet, are currently deployed in production AI coding assistants such as GitHub Copilot Chat [48]. Baseline performance results without adversarial perturbations are presented in Table II. On average, without any adversarial transformations, the models achieve a 60.79% pass rate, showcasing reasonable performance on these benchmarks and making them strong candidates for testing robustness against adversarial code transformations. We discuss the observed performance differences between our in-context task and original EvalPlus in Appendix E.
TABLE II: pass@1 performance of SoTA LLMs on popular code completion datasets.
Model HumanEval+ MBPP+
Claude 3.5 Sonnet v2 (2024/10/22) 70.73 67.29
GPT-4o 84.76 75.27
Artigenz Coder DS 6.7B 57.31 42.55
CodeQwen 1.5 7B Chat 68.29 67.02
DeepSeek Coder 6.7B Instruct 67.07 46.81
Llama 3 8B Instruct 32.32 50.00
For reproducibility, we set the sampling temperature to 0.0 to enable greedy decoding [^2] Anthropic API notes that setting temperature 0.0 does not guarantee complete determinism for its models.. For open-source models, we execute our attack with five random seeds. Due to resource constraints, we evaluate closed-source models with a single full run, supplemented by limited runs on small samples of each dataset using five distinct random seeds to measure variance (see Appendix F). Furthermore, to demonstrate our support for threat models that prohibit retrieving token probabilities (for example, at this time Anthropic API does not support returning token probabilities), we run our attack both with the token probabilities feedback, which allows us to use a perplexity based deep greedy composition, and without.
Results. Table III and Table IV present the attack performance and attack naturalness of GCGS.
@@
TABLE III: Performance of GCGS and GCGS+P (perplexity-guided) attacks on code generation using HumanEval+ and MBPP+. Results show mean ± std over 5 seeds for open-source models and single runs for closed-source models (limited 5-run analysis in Appendix F). Bold indicates best attack variant per model by ASR.

### v2

@@ 5 Greedy Cayley Graph Search
The goal of the XOXO attack is to modify the input code through semantics-preserving adversarial transformations that deceive the LLM, without changing the code’s underlying logic. Simple transformations include renaming variables or reordering independent statements. These transformations can change model output and confidence, as also shown by prior works [60, 28]. These simple, atomic transformations can be composed to create a vast space of potential transformations. The attack must explore this space to identify transformations that induce incorrect model outputs.
We observe that applying multiple confidence-reducing transformations often has a monotonic effect i.e. as more of these transformations are composed, the model’s confidence decreases consistently. We test this hypothesis in § F. Our search algorithm, GCGS, leverages this monotonicity to find a transformation composition that lowers model confidence below a threshold, resulting in incorrect output. We observe that the transformation space forms a tree structure, where each node (a composite transformation) is generated from a sequence of atomic transformations, and all nodes at a given level in the tree contain the same number of transformations. GCGS leverages the tree structure by representing it as a Cayley Graph [36], with the attack strategy executing a greedy walk through the graph to exploit the monotonic decrease in model confidence along the path.
@@ 5.1 Preliminaries
We consider a generating set $G$ of atomic transformations that generates the entire group of complex transformations. Each transformation $g_{i}\in G$ maps a code snippet $\mathcal{C}$ to $\mathcal{C}^{\prime}$ through atomic changes, such as replacing every occurrence of an identifier foo with bar, while preserving code semantics. For each transformation $g_{i}$, there exists an inverse transformation $g_{i}^{-1}\in G^{-1}$ that reverses its effect (e.g., replacing bar back to foo), such that their composition yields an identity transformation.
Since transformations in $G$ can be composed without restrictions, this set forms a free group $F(G)$, where each element represents a transformation sequence from $G\cup G^{-1}$. To systematically explore potential transformation sequences, we can represent this group using a Cayley Graph. For a free group, this graph becomes an infinite tree $\mathcal{T}$ as shown in Figure 3. In $\mathcal{T}$, each vertex represents an element of $F(G)$ (a composite transformation), and each edge represents the application of a single transformation $g\in(G\cup G^{-1})\setminus{e}$. Unlike other tree structures, Cayley graphs naturally handle cases where different transformation sequences when composed produce identical code snippets.
@@ 5.2 Problem Formulation
Consider a code model $\mathcal{M}:\ \mathcal{C}\rightarrow\mathcal{Y}$, mapping code snippets to an output space $\mathcal{Y}$ (e.g., class labels for classification tasks or token sequences for generation tasks). For many downstream tasks, even with black-box access to $\mathcal{M}$, we can approximately measure the model’s confidence in its predictions. Let $\alpha:\mathcal{C}\rightarrow[0,1]$ be a confidence scoring function. For classification tasks, $\alpha(c)$ can be derived from the probability distribution over classes [66, 71]. For generation tasks with current LLMs, we can compute $\alpha(c)$ using perplexity or prediction stability. This provides us with a continuous measure of the model’s certainty in its predictions, where lower values of $\alpha(g_{i}(c))$ indicate that applying transformation $g_{i}$ makes the model less confident about its output.
The model’s sensitivity to semantics-preserving transformations suggests it relies partly on spurious patterns rather than true understanding for its decision making. Consequently, applying multiple semantic-preserving changes can progressively degrade the model’s understanding of the program, manifesting as decreased confidence in its predictions. For transformations $g_{i},g_{j}\in G$, we expect their composition to result in lower confidence. We test this hypothesis in § F, where we find that composing transformations that individually decrease model confidence yields compounded confidence decreases, exhibiting monotonicity with respect to path length in $\mathcal{T}$. This monotonicity suggests a greedy strategy for finding adversarial transformations. Following paths of decreasing confidence in $\mathcal{T}$, we can efficiently reach composite transformations causing model mispredictions.
@@ 5.3 GCGS Algorithm
Leveraging the monotonicity property, GCGS finds a path to a transformation $\tilde{g}$ such that $\mathcal{M}(\tilde{g}(c))\neq\mathcal{M}(c)$. It explores the Cayley Graph $\mathcal{T}$ in two phases (Algorithm 1):
Shallow Exploration. GCGS begins by sampling a set $G^{R}\subset(G\cup G^{-1})\setminus{e}$ of $R$ generators. For each $g\in G^{R}$, it computes and stores the model confidence $\alpha(g(c))$ in a $g$-$\alpha$ map $A$. If any atomic transformation causes a model failure, the transformed code snippet is returned.
Algorithm 1 GCGS
@@
Deep Greedy Composition. If no atomic transformation succeeds, GCGS uses the stored confidence values to greedily compose transformations. Starting with the identity transformation $\tilde{g}=e$, it iteratively composes $\tilde{g}$ with generators from $G^{R}$, prioritized in order of increasing confidence values in $A$. This implements a guided descent through $\mathcal{T}$ towards likely failure points. Moreover, the inverse transformations in the generating set ($G^{-1}$) allow GCGS to revert any applied transformation along the greedy walk. GCGS repeats these two phases, maintaining the confidence map $A$ across iterations until it finds an adversarial example or reaches the query limit. GCGS implementation is detailed in § A.
5.4 GCGS with Warm-up
In the shallow exploration phase of the GCGS, randomly sampling from $(G\cup G^{-1})\setminus{e}$ to form $G^{R}$ can be query-inefficient as the sample may contain fewer confidence-reducing transformations. In practice, certain transformations might consistently be more effective at reducing model confidence across similar code snippets. We can exploit this pattern to make GCGS more efficient.
Consider an attacker with access to code snippets $C^{W}$ drawn from the target snippet distribution. We use $C^{W}$ in an offline stage to learn which transformations are most effective, warming up our attack to sample $G^{R}$ more intelligently during shallow exploration. We split $C^{W}$ into training set $C^{T}$ and validation set $C^{V}$. Over multiple rounds, we randomly sample $G^{R}$ from $(G\cup G^{-1})\setminus{e}$ and record $\alpha(g(c))$ for each $g\in G^{R}$ and $c\in C^{T}$. Using the average confidence drop of each transformation in $G^{R}$ on $C^{T}$, we run GCGS on $C^{V}$ to validate if the current sample of $G^{R}$ is better than the previous round. The warm-up procedure keeps refining the set $G^{R}$ until it either saturates, with GCGS’s performance on $C^{V}$ starting to deteriorate, or the maximum number of rounds is reached.
@@ 6 Evaluation
We evaluate the efficacy of GCGS at attacking models performing code generation and reasoning. First, in § 6.1, we devise an in-context code generation task simulating how AI assistants will formulate code generation prompts with supplementary context. This allows us to assess how vulnerable the underlying LLMs are to context poisoning through semantics-preserving code transformations. Second, we evaluate GCGS on the CWEval dataset, a benchmark designed to assess the security of LLM-generated code, in § 6.2. Finally, in § 6.3, we compare GCGS against SoTA adversarial attacks on two security-critical code reasoning tasks: defect and clone detection.
@@
Model Confidence (Generation vs Reasoning) In code reasoning tasks involving classification, adversarial attacks can use the model’s output probabilities as an optimization signal. For code models with this guidance, we compare GCGS and GCGS with warm-up (GCGS+W) against SoTA black-box adversarial input generation methods which use model confidence to guide their exploration. However, existing approaches fail when attacking models on code generation (detailed in § C.1). Consequently, in the generation setting, we evaluate GCGS both with perplexity feedback from generated inputs (GCGS+P) and without it when token probabilities are unavailable (GCGS).
@@
Evaluation Metrics. The performance of our attack is measured using two metrics: (i) Attack Success Rate (ASR) is the percentage of cases where an attack transforms correct model outputs into incorrect ones. This applies to classification (model’s prediction changes from correct to incorrect class) and code generation (generated code changes from passing to failing test cases), and (ii) Number of Queries (# Queries) refers to the mean number of model queries per attack, indicating the attack’s efficiency under real-world constraints like rate limits and cost. In addition, we evaluated the quality and naturalness of the adversarial examples generated by GCGS (detailed in § C.4).
6.1 In-Context Code Generation
@@
Task Description. We formulate a new task to simulate a code generation task with benign, AI assistant-supplied context. For code generation, we use the industry-standard HumanEval+ (164 problems) and MBPP+ (378 problems) datasets from EvalPlus [40]. Both datasets contain Python functions with natural language descriptions and input-output examples, with performance measured using the pass@1 metric – percentage of correct solutions a model generates on its first attempt.
@@
To simulate a realistic code generation scenario with an AI assistant, we augment each target problem with three randomly sampled, solved examples from the same dataset. The prompt instructs the model to generate a solution for the target problem while adhering to the coding style and naming conventions of the provided context examples. The full prompt template is detailed in § D.
r
@@
Table 1: Performance of GCGS and GCGS+P (perplexity-guided) attacks on code generation (HumanEval+ and MBPP+) and vulnerability injection (CWEval/Python). Results show mean ± std over 5 seeds for open-source models and single runs for closed-source models (limited 5-run analysis in § H). Bold indicates best attack variant per model by ASR.